MUI-FReq

Fido File Request utility Version 1.1 29/04/1994

by Sandro Tolaini

Permission is granted to make and distribute verbatim copies of this manual, and to copy and distribute translations into another language, provided the copyright notice and this permission notice are preserved on all copies.

1 What is MUI-FReq

Well... if you are a Fido Point, you’ll guess what is MUI-FReq. However, that’s it...

MUI-FReq is a tool to create and manipulate the ‘.REQ’ files used by TrapDoor and other Fidonet mailers to request files from nodes. It has a lot of features not found on others similar programs. They are:

• An Easy-To-Use MUI interface. MUI is copyrighted by Stefan Stuntz, and is not included in the distribution archive. You can obtain it from various BBS, or directly from Stefan.

• The capability to create and use backups for the filelists. Never happened you call your favourite BBS to get the latest 100 PD utilities, and some of them aren’t available? You have to rewrite the 100 names, and hope the next time they are there...

• An Outbound scan facility that allows you to see what nodes have pending file requests. So you can load them and poll the node.

• An Inbound scan facility that allows you to delete from the filelist the files that you have just downloaded. This is great used with the backup option.

• Full localization of the program. Only the Italian catalog is available, but you can do your own translation and send it to me. It will be included in the next release.

• Direct polling from the program, using your favourite mailer. The address of the node can be put in the command string, making multi-polling very easy.

• A filelist browser, useful for selecting filenames directly from a BBS file list. This is not yet implemented, but will be in the next release.

• An ARexx interface, to use MUI-FReq with your favourite reader. Also this isn’t yet implemented, but will be very soon.

Have fun with MUI-FReq!

2 Distribution

MUI-FReq is distribuited under the concept of ShareWare: if you use this program, you are required to send 10.000 Lit. or the equivalent of 8 US$ in any currency to the author. This program is not crippled in any way, however future releases depends heavily on how much feedback it will have.

If you can’t afford paying money for MUI-FReq, send me at least a postcard with your comments. For the address, see Chapter 3 [Author Info], page 2.

3 Author Info

You can reach me at the following addresses:

  Sandro Tolaini
  Via Giannone 3
  50047 Prato (FI) ITALY

or with E-Mail:

  2:332/113.3@fidonet
  39:102/507.3@amiganet

If you want to have a reply, use only E-Mail. Thanks.

4 Requirements and installation

MUI-FReq requires Kickstart and Workbench 2.04 (WB 2.1 if you want to use localization) or higher to run. Also, you need MUI 2.0, which is available in many BBS.

To install MUI-FReq do the following:

• Drag the MUI-FReq icon to the drawer you want to install the program on.

• If you want to use localization, copy the ‘Catalogs’ directory to your ‘LOCALE:Catalogs’ directory.

This is all you have to do... However, in the next release I will include an Installer script, for fast and easy installation.

5 Main window

The main window consists of three areas, from the top to the bottom.

• Just below the window title, you can see a message area, where MUI-FReq displays errors or other messages for the user.

• Below, you have the filelist and the related gadgets, that lets you manage the file names to be requested. Gadgets description follows: Fido address lets you to specify the fidonet node you want to edit the filelist for. The address must be in the form z:n/f.p, where z, n, f and p are decimal numbers. To allow MUI-FReq recognize the address change, press RETURN when you have edited the address. Do not use TAB or the mouse, or the address won’t be accepted. The popup gadget near the address will scan the Outbound directory for pending filerequests. Each node that has a ‘.REQ’ file (or a backup) will be displayed. Double-clicking on an entry will change the current address and load the proper filelist. If you start MUI-FReq from CLI, you can insert in the command line a default value for this gadget. For example, MUI-FReq 2:332/113.0 will start the program with 2:332/113.0 as the default node. This is expecially useful if you substitute MUI-FReq to one of the built-in filerequest tools of fidonet editors like Mail Manager or Spot. Add will add a new filename to the list, and activate the filename string gadget, to allow you editing of the name. Press RETURN to insert the filename, and MUI-FReq will be ready to accept another filename. When you have finished, just press RETURN on the empty filename string. Remove will delete the selected entry from the filelist. This can also be done by selecting a file, clearing the filename string gadget with Amiga-X and pressing RETURN. Up will move up the selected entry. Down will move down the selected entry. Sort will sort the filename list alphabetically. Clear will clear the whole filelist.

• Then, you have the action gadgets, that allows you to do various useful things:

Load list will load (if exists) the filelist for the current address. A proper message will be displayed in the message area. Save list will save the current filelist in the form of a ‘.REQ’ file for the current address. If the current filelist is empty, the ‘.REQ’ file will bel deleted. Poll will execute the configured poll command, in order to call the current node. The filelist will be saved before polling.

Node Info will pop up a requester with some useful informations about the current node. This function is available only if you have installed the traplist.library which is distributed with the TrapDoor mailer. Scan Inbound will check for every name in the current filelist if there is the corresponding file in the Inbound directory. If a file is found, it will be removed from the filelist. Very useful! Configuration... will open the configuration panel. See Chapter 6 [Configuration window], page 4. About... will show you some useful informations about the author.

6 Configuration window

The configuration window lets you specify the paths and the options for MUI-FReq. It is divided into two sections, explained below.

• The paths and options gadgets. They are: Outbound path is the same path configured in your mailer for the outbound files. It is where files to be sent are kept. Inbound path is the same path configured in your mailer for the inbound files. It is where files received are kept. Poll command is the command string to be executed in order to call your mailer. If you place a %s in the command string, it will be substituted by the Fidonet address of the node to be called, in the form z:n/f.p. Use filelist backups allows you to save and load backups for the filelists. They are saved as ‘.BAK’ files in the outbound directory. If a backup list is loaded (when the ‘.REQ’ file for a node is not found) a proper message will be displayed. Scan inbound after poll will force MUI-FReq to scan the inbound directory for received files just after polling a node. Save list before quitting will cause the program to save the current filelist before quitting. Normally, you have to do it by hand. Convert filenames to uppercase will cause every name inserted in the filelist to be converted in uppercase. Some Fidonet file requester server need an uppercase name to work properly (especially those running on MS-DOS systems).

• The actions gadgets. They are:

Save will save the configuration, so it will be loaded at program startup. Use will use this configuration, but when you quit the program it will be lost.

7 Acknowledgments

My thanks go to the following people:

• Debby, for his smile that changed my life.

• Gianmarco Giovannelli, for his suggestions and beta testing.

• All the people that will register this program, for their support.

• Amiga, for being the best computer!